# Float Self-Tagging

### v2

@@ 4. Improved Predictability and Coverage
But which floats are most common in programs? To answer this question the float-intensive programs of the R7RS benchmark suite are instrumented to create a profile of all the floats computed by those programs. The results are shown in Figure 6. The table contains 32 rows, one for each combination of the 5 most significant bits of the float’s 11-bit exponent field, which is indicated in the first column. For now ignore the second column. Each of the right columns indicates what proportion of computed floats have a specific combination of most significant exponent bits. The computed floats are clearly clustered around 2, specifically in the range $1.1\times 10^{-19}$. $3.7\times 10^{19}$, and the value 0.0 also occurs frequently for some programs. It is clear that other scientific programs may exhibit profiles that use other ranges of floats. In particular, Infinity and NaN may show up more often in some types of computations. It is apparent that many programs will operate with floats that are predominantly those around 2 on a logarithmic scale, and also 0.0, Infinity, and NaN.
@@ 4.1. Self-Tagging with 3 Tags and 4 Tags
The float profile of Figure 6 has the nice property that all floats computed have the 3 most significant bits of the exponent field that are either 000 (green rows), 011 (blue rows), or 100 (orange rows). Moreover, among the values captured by 000, only $\pm$0.0 are used.
These 3 exponent bits can be shifted using a bit rotation to align them with the lower 3 tag bits as shown in Figure 7. Thus a float is encoded to a tagged value by a 4-bit rotation to the left (or 60-bit to the right). A tagged value is decoded to a float with the inverse rotation.
Contrarily to the implementation of self-tagging in Section 3, these tags capture intervals of contiguous values. This contiguity within intervals prevents small variations from causing switches from a self-tagged value to a heap-allocated float.
Test programs did not compute the float values $\pm$±Infinity and NaN, but those values would be assigned the tag 111 (gray rows). Assigning this fourth tag to self-tagged floats might be useful to avoid heap-allocated floats in programs frequently computing these values.
Given the expected rarity of the remaining floats there isn’t much of an advantage to represent them using tagged pointers. Instead a representation using a generic pointer with a header type is a better tradeoff as it allows using the tag that would have been used for some other, more frequently used type.
4.2. Self-Tagging with 2 Tags
As is shown in the float profile of Figure 6 the tag 000 is only used to capture the values $\pm$0.0. It is possible to avoid using the tag 000 for self-tagging by preallocating the values $\pm$0.0 using the generic pointer representation. When a float must be encoded as an object it is compared to zero, in which case one of the preallocated zeroes is returned, thus avoiding a new allocation. The same can be done for the values $\pm$±Infinity and NaN, which are expected to be the most frequent values tagged with 111. If this is done then only 2 tags are needed to cover the most frequent cases: 011 and 100.
Unfortunately, this approach may raise the number of failed branch predictions due to the test for the special values $\pm$0.0, $\pm$±Infinity, and NaN. Consequently it may be a more performant tradeoff to avoid testing for some of the special values which are instead handled with a heap-allocation. The best choice will depend on the context, for example the float profile of Figure 6 would suggest not testing for $\pm$±Infinity and NaN.
@@ 4.3. Self-Tagging with 1 Tag
Self-tagging can also be achieved with a single tag by a simple transformation of the most significant exponent bits. The 5 most significant bits of the exponent are added to 1 and the middle 3 bits are used as the tag (using a rotation 5 places to the left). This corresponds to the second column of Figure 6. Note that the tag 000 (bold pink rows) covers all the useful ranges, including $\pm$±Infinity and NaN that are not covered by the 2-tag and 3-tag variants.
@@
This variant of self-tagging has the advantages of covering all the useful ranges of float values and using a single tag, which leaves more tags available for representing other objects. The only downside when compared to the 4-tag variant is that it covers proportionately fewer ranges of floats, so there is a risk of more heap-allocated floats.
@@ 4.4. Self-Tagging with a Different Exponent Bias
A small change in the exponent bias of the 1-tag variant can be used to get another self-tagging variant. If $2\times tag_{1}$ is added to the highest 5 bits and the middle 3 bits are examined, then when they are equal to $tag_{1}$ or $tag_{2}=tag_{1}-1$ they are self-tagged floats. This effectively doubles coverage by including both the light and bold pink rows of the second column of Figure 6. In other words the following (absolute value) ranges will be self-tagged: $0..3.8\times 10^{-270}\cup 5.9\times 10^{-39}..6.8\times 10^{38}\cup 1.1\times 10^{270}..Infinity/NaN$. To put this into perspective, this is a superset of the IEEE754 32-bit floats, so any floating point value that can be represented as a IEEE754 32-bit float can be self-tagged with this variant which uses 2 of the 8 3-bit tags. This is very likely a better use of 2 tags as the 2-tag variant of Section 4.2 does not cover the ranges containing 0, Infinity, and NaN which need special treatment that slows down execution.
When compared to the 1-tag variant, there will be no difference in heap allocations for programs with similar float profiles as Figure 6. However, as will be explained in the next section, it may yield more compact and faster code, in essence because the bias can be added after the rotation, and the need for a bias disappears when $tag_{1}$=000.
@@ 5. Implementation
This section discusses the implementation of the float-related operations of the dynamic language. The types f64 and f32 will refer to IEEE754 64-bit and 32-bit doubles and the types i64 and i32 will refer to 64-bit and 32-bit integers. Moreover, the type object will refer to the values of the language (the sum type of floats, integers, booleans, and other objects). To better understand the low-level implications of the implementation and give a sense of the execution time performance, assembly code for the Intel x86 architecture is shown to illustrate the details of the implementation on that popular architecture. Although this section focusses on a 64-bit implementation where the type object is an i64 and floats are f64, it is fairly straightforward to implement the operations similarly on a 32-bit implementation where the type object is an i32 and floats are f32.
@@ 5.1. Single Tag Testing
Note that by using al only the lower 8 bits of rax are accessed, leading to smaller constants and instruction encodings. However, this sequence modifies rax so if the value is needed after the tag test, as often will be the case, then an additional instruction and register are required to keep a copy. In the special case of tag=0, the two first instructions can be replaced with “test al, 7”, which is a bitwise-and that tests if all 3 lowest bits of rax are zero without modifying rax. In the general case when $\neq$0 the test instruction can be used by combining it with a lea (Load Effective Address) instruction that copies to a temporary register the source register plus a constant that cancels tag (modulo 8):
lea ebx, [eax+(8-$tag$)]
@@ 5.2. Multiple Tag Testing
When it is necessary to check if an object’s tag is one of a set of tags it is often possible to do this more efficiently than a sequence of single tag checks. This is particularly the case when there is leeway in the assignments of tags as is often the case.
[…]
lea ebx, [eax+(8-3)] # 011 -> 000 and 100 -> 001
test bl, 6
jz tag_matches_011_or_100
This technique can be extended when testing the 4 tags 000, 011, 100, and 111 of the 4-tag self-tagging approach:
lea ebx, [eax+(8-7)] # 000 -> 001, 011 -> 100, 100 -> 101, and 111 -> 000
test bl, 2
jz tag_matches_000_or_011_or_100_or_111
[…]
@@
The bit index in ax is obtained mod 16, so the byte 0b00011001=0x19 (all 0’s except at bit indices 0, 3, and 4) is repeated twice in register bx. Using the bt instruction has the advantage of neither modifying the rax register nor the register holding the set of tags, so the cost of the mov instruction can be amortized over multiple checks of that set of tags (a possibly near zero cost if that register is globally reserved).
On architectures without a bit test instruction, a dynamic count shift of the tag set register could be used to achieve the bit indexing. If this is done by modifying the tag set register the initialization can’t be shared by multiple tag tests. On most 3-address RISC architectures the shift can be done non-destructively. On ARM A64, here is a short instruction sequence that tests the 3 low bits of 64-bit register x1 for a tag of 0, 3, or 4:
@@
The cost of the check does not depend on the number of tags tested. This can be advantageous if a tagged pointer is used for heap-allocated floats, say with tag h. In that case, a check for a float (either self-tagged or heap-allocated) can be done by adding h to the tag set. Once it is known that an object is a float, it is easy to check for the single tag h to discriminate between the self-tagged and heap-allocated representations.
@@
To convert a f64 to an object, it must first be converted to an i64, and then a constant $bias$ must be added modulo $2^{64}$ followed by a bit-rotation by $r$ places left. For the 1-tag variant, $bias=(1+2\times tag)\times 2^{58}$ and $r=5$. For the 2-tag, 3-tag, and 4-tag variants, $bias=0$ (no bias) and $r=4$. After this there is a check to verify if the resulting object’s tag is in the set of self-tagged tags. If this is the case the conversion is done, otherwise an out of line routine can be called to heap-allocate a float and return an appropriately tagged pointer. As an example, if the f64 value is in the 64-bit float register xmm0, then the following x86 code will set register rax to the corresponding object when the 3-tag self-tagging variant is used:
@@
The following x86 code is appropriate for the 1-tag self-tagging variant using 000 as the tag:
@@
For the variant explained in Section 4.4 which uses 2 tags with no special handling of zeroes, it is more efficient to do the rotation before adding the bias because then the bias is the small constant $tag_{1}$, and in the special case of $tag_{1}$=000 the bias vanishes. For example, here is the code when $tag_{1}$=011:
@@ 5.5. C Implementation
This section explains how the above operations can be implemented in C, as this is a common implementation language, in particular it is used by the Bigloo and Gambit Scheme to C compiler used in the experiments of Sections 3 and 6.
@@
Using this pure C definition with GCC version 13.2.0 and clang version 18.1.3 on x86-64 results in both cases in the non-destructive approach based on the bit test instruction. The expression n & 31 is optimized out by those compilers because the shift instruction on x86 implicitly does this operation. On older versions of those compilers that would implement this with a destructive shift instruction, the bit test approach can be achieved using the following definition that uses an asm statement:
@@
6. Experiments of Self-Tagging using the Exponent Bits
The experimental setup from Section 3 is replicated to test float self-tagging using high bits of the exponent. Three variants are compared to the original version of Bigloo that implements standard tagged objects.
@@ 6. Experiments of Self-Tagging using the Exponent Bits
  • 3-tag self-tagging uses the tags 000, 011 and 100, thus self-tagging floats in the green, blue, and orange ranges from Figure 6. The remaining floats are represented by generic pointers with a header type, and thus are heap-allocated.
@@
  • 2-tag self-tagging tests the impact of expending the tag 000 for self-tagging. It only uses the tags 011 and 100 for float self-tagging and reclaims the tag 001 to represent all remaining floats as tagged pointers. The floats $\pm$0.0 are preallocated and the float boxing operation contains an explicit zero check to return the preallocated zeros.
  • 1-tag self-tagging tests the 1-tag variant where the tag 001 is used for self-tagged floats.
Figure 8 shows the memory allocations of these variants. All three variants allocate fewer floats on all float-intensive benchmarks. In fact, all variants have been verified to allocate no float at all (see Figure 6) with the only exceptions being the preallocated $\pm$0.0 by the 2-tag variant. The remaining heap usage is for non-float objects. Similarly to the previous experiment (Figure 5(a)), memory usage of other benchmarks remains mostly unaffected.
Figure 8. Heap allocation comparison between the three variants of self-tagging on exponent bits. The 3-tag variant uses the tags 000, 011 and 100. The 2-tag variant uses the tags 011 and 100, but preallocates 0.0 and -0.0. The 1-tag variant uses 3 bits of the biased exponent (see Figure 6). The y-axis shows heap allocations relative to the original Bigloo version, which uses tagged pointers for all floats. 1.0 indicates no change, while 0.0 indicates that all allocations were removed.
Figure 9 shows faster execution times for all three variants of self-tagging. All float-intensive benchmarks execute faster than the original Bigloo implementation, with the exception of fft on the 2-tag variant. In particular, the mbrot and nucleic benchmarks are both faster, confirming that self-tagging on exponent bits eliminates the branch misprediction penalty incurred by self-tagging on mantissa low bits (see Section 3). Despite small individual variations, other benchmarks remain generally unaffected.
On the float-intensive benchmarks, execution time with the 3-tag variant, which makes use of the tag 000, is on average 0.41$\times$ that of the original Bigloo version, slightly more than two times faster. This is significantly lower than the 2-tag variant (0.48$\times$) which avoids the tag 000. In particular, the benchmark fft, which predominantly uses the values $\pm$0.0, is slower in the 2-tag variant despite zeros being preallocated. The 1-tag variant (0.42$\times$) is similar to 3-tag on float benchmarks, and both variants display no performance change on non-float benchmarks (1.01$\times$ for 1-tag, and 1.03$\times$ for 3-tag). Due to the 1-tag variant not using 000, and leaving more tags available, it is a more versatile candidate and is thus used as the default variant for the rest of this paper.
Figure 9. Execution time comparison between the three variants of self-tagging on exponent bits. The 3-tag variant uses the tags 000, 011 and 100. The 2-tag variant uses the tags 011 and 100, but preallocates 0.0 and -0.0. The 1-tag variant uses 3 bits of the biased exponent (see Figure 6). The y-axis shows execution time relative to the original version. 1.0 indicates no change, lower means faster execution than the original Bigloo version, which uses tagged pointers.
6.1. Effect on Garbage Collection
Execution time (s)
fft Refer to caption
fibfp Refer to caption
mbrot Refer to caption
nucleic Refer to caption
pnpoly* Refer to caption
ray Refer to caption
simplex Refer to caption
sum1 Refer to caption
sumfp Refer to caption
Preallocated Heap Data Size
Figure 10. Execution time of float-intensive R7RS benchmarks with extra live data allocated on heap before execution. Benchmarks are executed in unsafe mode (without typechecks) to specifically gauge the impact on garbage collection. The x-axis shows the size of preallocate data in bytes on a logarithmic scale. An extra point is added at $x=0$ for execution time without preallocated data. The y-axis represents execution time in seconds. Dashed lines represent systems that do not implement self-tagging (original Bigloo, original Gambit, and Chez Scheme). Solid lines are Bigloo and Gambit with self-tagging (1-tag).
*pnpoly’s execution time with Bigloo (orig) is about 230 seconds, which is cropped out for readability.
Float self-tagging improves performance by allocating fewer floats on the heap, thus relieving pressure on the garbage collector. Hence, the impact of self-tagging is more pronounced for programs that spend more time doing memory management tasks. This section presents an experiment to measure execution time on float-intensive R7RS benchmarks when the heap contains live data, thus putting more pressure on the garbage collector.
The R7RS benchmarks suite is modified to allocate a vector before the execution of each benchmark. The size of the allocated vector is given as a parameter of the benchmark. Each benchmark is then executed normally with vectors of size ranging from $10^{4}$ to $10^{8}$, which correspond to about 80 kB and 800 MB of memory respectively. An execution is also done with no preallocated vector. To specifically measure the impact of self-tagging on garbage collection, benchmarks are executed in unsafe mode, which allows unsafe optimizations such as the removal of safety typechecks.
For this experiment the float self-tagging approach has been implemented in the Gambit 4.9.6 (Marc Feeley, 2025) optimizing Scheme compiler. The benchmarks are executed with Bigloo and Gambit, with and without self-tagging (1-tag), and also the Chez Scheme 10.1 (R. Kent Dybvig, 2025) optimizing Scheme compiler without self-tagging. These compilers are chosen due to their status as state-of-the-art Scheme implementations and the fact they employ distinct garbage collection techniques. Chez Scheme uses a generational garbage collector. Gambit uses a stop-and-copy garbage collector with bump allocation. Bigloo uses the Boehm-Demers-Weiser Garbage Collector (Boehm and Weiser, 1988).
Figure 10 compares execution times of these Scheme implementations. With a nearly empty heap, and thus little strain on the garbage collector, the bump allocator of Gambit without self-tagging generally outperforms all other implementations since it does not incur the cost, albeit small, of encoding and decoding self-tagged floats. However, once the heap is preloaded with 1 to 10 MB (which is roughly the size of the data cache) the strain on all garbage collectors starts to outweigh the cost of encoding and decoding self-tagged floats, thus Bigloo and Gambit with self-tagging hereafter outperform all implementations without self-tagging.
The threshold of 1 to 10 MB of preallocated data beyond which self-tagging increasingly outperforms other implementations suggests that experiments in other sections are in fact worst-case scenarios. On real life programs, where it is reasonable to expect some persistent data in the heap, self-tagging can be expected to offer a more substantial improvement.
6.2. Comparison to NaN-Boxing and NuN-Boxing
Bigloo also contains implementations of NaN-boxing and NuN-boxing which can be enabled through a configuration option.
Figure 11 compares Bigloo’s original representation (tagged pointers) to NaN-boxing, NuN-boxing, and 1-tag float self-tagging. On float-intensive benchmarks, the average execution time with self-tagging (0.42$\times$ of original version) is higher than with NaN-boxing (0.30$\times$) and NuN-boxing (0.34$\times$). On non-float benchmarks the average relative execution time of self-tagging (1.01$\times$), NaN-boxing (1.03$\times$), NuN-boxing (1.03$\times$) all remain similar to the original version. However, on individual benchmarks the performance of self-tagging remains generally unchanged while significant variations are observed with NaN/NuN-boxing.

### v3

@@ 2.1. Float Coverage
The results are shown in Figure 5. The table contains 32 rows, one for each combination of the 5 most significant bits of the float’s 11-bit exponent field, which is indicated in the first column (for now, ignore the second column). Each of the right columns indicates what proportion of computed floats have a specific combination of most significant exponent bits. The computed floats are clearly clustered around 1.01.01.0, specifically in the range $1.1\times 10^{-19}$. $3.7\times 10^{19}$, and the value 0.0 also occurs frequently for some programs. In fact, R7RS benchmarks have a distribution of floats that appears even more tightly concentrated near 1.01.01.0 than that reported by (Zurstrassen et al., 2023, Fig. 7), shown in the rightmost column. Hence, other scientific programs may exhibit profiles that use slightly wider ranges of floats. It is apparent that many programs will operate with floats that are predominantly those around 1.01.01.0 on a logarithmic scale, and also 0.0.
@@ 2.2. Self-Tagging with 3 Tags and 4 Tags
The float profile of Figure 5 has the property that almost all floats computed have the 3 most significant bits of the exponent field that are either 000 (green rows), 011 (dark blue rows), or 100 (light blue rows). Moreover, among the values captured by 000, only $\pm$0.0 are used.
These 3 exponent bits can be shifted using a bit rotation to align them with the lower 3 tag bits as shown in Figure 4. Thus a float is encoded to a tagged value by a 4-bit rotation to the left (or 60-bit to the right). A tagged value is decoded to a float with the inverse rotation. This corresponds to the encoding discussed in the previous section (Figure 4).
Test programs did not compute the float values $\pm$±Infinity and NaN, but those values would be assigned the tag 111 (gray rows). Assigning this fourth tag to self-tagged floats might be useful to avoid heap allocated floats in programs frequently computing these values.
2.3. Self-Tagging with 2 Tags and Preallocated Zeros
As is shown in the float profile of Figure 5 the tag 000 is only used to capture the values $\pm$0.0. It can be freed for other uses by preallocating the values $\pm$0.0 using tagged pointer or generic pointer representation. When a float must be encoded as an object it is compared to zero, in which case one of the preallocated zeroes is returned, thus avoiding a new allocation. If this is done then only 2 tags are needed to cover the most frequent cases: 011 and 100. Unfortunately, this approach may raise the number of failed branch predictions due to the test for $\pm$0.0 (this is further discussed in Section 4.6).
@@ 2.4. Self-Tagging with 1 Tag
Self-tagging can also be achieved with a single tag by a different transformation of the most significant exponent bits. The 5 most significant bits of the exponent are added to 1 and the middle 3 bits are used as the tag (using a rotation 5 places to the left). This corresponds to the second column of Figure 5. Note that the tag 000 (bold pink rows) covers all the useful ranges, including $\pm$±Infinity and NaN that are not covered by the 2-tag and 3-tag variants.
@@
This variant has the advantages of covering all the useful ranges of floats computed by R7RS benchmarks and about 90% of floats reported by (Zurstrassen et al., 2023), while using a single tag, which leaves more tags available for other types.
@@ 2.5. Self-Tagging with a Different Exponent Bias
A small change in the exponent bias of the 1-tag variant can be used to get another self-tagging variant. If $2\times tag_{1}$ is added to the highest 5 bits and the middle 3 bits are examined, then when they are equal to $tag_{1}$ or $tag_{2}=tag_{1}-1$ they are self-tagged floats. This effectively doubles coverage by including both the light pink and bold pink rows of the second column of Figure 5. The following ranges will be self-tagged: $0..3.8\times 10^{-270}\cup 5.9\times 10^{-39}..6.8\times 10^{38}\cup 1.1\times 10^{270}..Infinity/NaN$. To put this into perspective, this is a superset of the IEEE754 32-bit floats, so any floating point value that can be represented as a IEEE754 32-bit float can be self-tagged with this variant which uses 2 of the 8 3-bit tags. This is likely a better use of 2 tags than the 2-tag variant of Section 2.3, which does not cover the ranges containing 0, Infinity, and NaN.
When compared to the 1-tag variant, there will be almost no difference in heap allocations for programs with similar float profiles as Figure 5. However, as will be explained in the next section, it yields more compact code, because the bias can be added after the rotation, and the need for a bias disappears when $tag_{1}$=000.
@@ 3. Implementation
This section discusses implementation details for float-related operations in dynamic languages. The types f64 and f32 will refer to IEEE754 64-bit and 32-bit doubles and the types i64 and i32 will refer to 64-bit and 32-bit integers. Moreover, the type object will refer to values in the language (the sum type of floats, integers, booleans, and other objects). To illustrate the low-level implications of the implementation and give a sense of the execution time performance, assembly code for the Intel x86 architecture is provided. This section focusses on a 64-bit implementation where object is an i64 and floats are f64, but it is straightforward to implement the operations similarly on a 32-bit implementation where the type object is an i32 and floats are f32.
@@ 3.1. Single Tag Testing
Note that by using al only the lower 8 bits of rax are accessed, leading to smaller constants and instruction encodings. However, this sequence modifies rax so if the value is needed after the tag test, as often will be the case, then an additional instruction and register are required to keep a copy. In the special case of tag=0, the two first instructions can be replaced with “test al, 7”, which is a bitwise-and that tests if all 3 lowest bits of rax are zero without modifying rax. In the general case when $\neq$0, the test instruction can be combined with a lea (Load Effective Address) instruction to copy the source register, plus a constant to cancel tag (mod 8), to a temporary register:
[…]
@@ 3.2. Multiple Tag Testing
When it is necessary to check if an object’s tag is one of a set of tags it is often possible to use a single branch instruction rather than a sequence of single tag checks. This is particularly the case when the implementer has leeway in the assignments of tags, as is often the case.
Checking that $tag\in\{tag_{1},tag_{2}\}$ where $tag_{2}=tag_{1}+2^{n}$ (mod 8) can be done just as efficiently as the single tag test by using the lea instruction to map $tag_{1}$ to 0 and $tag_{2}$ to $2^{n}$ and then using a mask of $7-2^{n}$ in the test instruction to ignore the bit that may be 0 or 1. For example, for 2-tag self-tagging (Section 2.5) with the tags 010 and 110 the check is:
lea ebx, [eax+(8-2)] # 010 -> 000 and 110 -> 100
test bl, 3 # 3 = 7-4
jz tag_matches_010_or_110
4-tag self-tagging (Section 2.2) with the tags 010, 011, 110, and 111 is even simpler:
test al, 2
jnz tag_matches_010_or_011_or_110_or_111
The general case of checking that $tag$ is in a set of tags can be done efficiently with the x86 bit-test instruction “bt n, i” that reads the bit at index i of n and puts it in the carry flag. The bt instruction comes in 16, 32, and 64 bit variants, but not 8 bit. For example, for 3-tag self-tagging (Section 2.2) with the tags 000, 011, and 100 the check is:
@@
The bit index in ax is obtained mod 16, so the byte 0b00011001=0x19 (all 0’s except at bit indices 0, 3, and 4) is repeated twice in register bx. Using the bt instruction has the advantage of neither modifying the rax register nor the register holding the set of tags, amortizing the cost of mov over multiple checks of that set of tags (a possibly near zero cost if that register is globally reserved).
On architectures without a bit test instruction, a dynamic count shift of the tag set register can achieve the bit indexing. If done by modifying the tag set register, the initialization can’t be shared by multiple tag tests. On most 3-address RISC architectures the shift can be done non-destructively. On ARM A64, this instruction sequence tests the 3 low bits of 64-bit register x1 for a tags 0, 3, or 4:
@@
The cost of the check does not depend on the number of tags tested. This can be advantageous if a tagged pointer is used for heap allocated floats, say with tag h. In that case, a check for a float (either self-tagged or heap allocated) can be done by adding h to the tag set. Once it is known that an object is a float, it is easy to check for the single tag h to discriminate between the self-tagged and heap allocated representations.
@@
To convert a f64 to an object, it is converted to an i64, added to a constant $bias$ modulo $2^{64}$, and bit-rotated left by $r$. For the 1-tag variant, $bias=(1+2\times tag)\times 2^{58}$ and $r=5$. For the 2-tag, 3-tag, and 4-tag variants, $bias=0$ (no bias) and $r=4$. After this there is a check to verify if the resulting object’s tag is in the set of self-tagged tags. If this is the case the conversion is done, otherwise an out of line routine can be called to heap allocate a float and return an appropriately tagged pointer. As an example, if the f64 value is in the 64-bit float register xmm0, then the following x86 code will set register rax to the corresponding object when 3-tag self-tagging is used:
@@
The following x86 code is appropriate for 1-tag self-tagging using 000 as the tag:
@@
For the 2-tag variant with no special handling of zeroes (Section 2.5), the rotation can be done before adding the bias to use a compact addition instruction (because then the bias is the small constant $tag_{1}$). In the special case of $tag_{1}$=000 the bias vanishes. For example, here is the code when $tag_{1}$=011:
@@
In summary, the self-tagging operations require few machine instructions but there are small variations depending on the instruction set and also the assignment of tags which can often be optimized for self-tagging without impacting the speed of operations on objects or small integers.
@@ 3.5. C Implementation
This section explains how the above operations can be implemented in C, as this is a common implementation language, in particular it is used by the Bigloo and Gambit Scheme to C compilers used in the experiments of Section 4.
@@
Using this pure C definition with GCC version 13.2.0 and clang version 18.1.3 on x86-64 results in the non-destructive approach based on the bit test instruction. The expression n & 31 is optimized out by those compilers because the bit test instruction on x86 implicitly does this operation. On older versions of those compilers that would implement this with a destructive shift instruction, the bit test approach can be achieved using the following definition that uses an asm statement:
@@
4. Experiments
This section evaluates float self-tagging and demonstrates that it presents an interesting performance trade-off between programs that perform frequent operations on objects but few floating-point operations, and float-intensive programs where this balance is reversed. This evaluation is conducted through experiments that measure performance metrics across different object representations (both self-tagging and existing ones), multiple machine architectures, and two state-of-the-art Scheme implementations: the Bigloo (commit 5b1118(Manuel Serrano, 2024) and Gambit v4.9.7 (commit 768900(Marc Feeley, 2025) Scheme compilers. Using two independently developed compilers helps to demonstrate that self-tagging can be adapted to preserve the different design philosophies and historical implementation choices of these systems.
The following self-tagging variants are evaluated:
@@ 4. Experiments
  • 3-tag self-tagging uses 3 tags for self-tagging floats corresponding to the green, dark blue, and light blue ranges from Figure 5. The remaining floats are heap allocated. Bigloo uses the tags 000, 011 and 100, and heap allocated floats are represented with generic pointers with a header type. Gambit offsets the tags by 3 (i.e. it uses 011, 110 and 111) so that the tag 000 is kept for tagging small integers, and tagged pointers with the tag 010 are used for the heap allocated floats.
  • 4-tag self-tagging (Gambit only) is like the 3-tag variant but uses the tag 010 to self-tag floats in the gray range from Figure 5. Allocated floats are represented with generic pointers.
@@
  • 2-tag self-tagging with preallocated zeros (Bigloo only) tests the impact of expending the tag 000 for self-tagging. It only uses the tags 011 and 100 for float self-tagging the dark blue and light blue ranges from Figure 5 and reclaims the tag 001 to represent all heap allocated floats as tagged pointers. The floats $\pm$0.0 are preallocated and the float boxing operation contains an explicit zero check to return the preallocated zeros.
  • 2-tag self-tagging (Gambit only) tests the variant that uses the tags 010 and 110 for float self-tagging the bold pink and light pink ranges from Figure 5.
  • 1-tag self-tagging tests the variant that uses a single tag for self-tagging the bold pink ranges from Figure 5. Bigloo uses the tag 001 and Gambit uses the tag 110. heap allocated floats are represented with tagged pointers.
Additionally both compilers can be configured to represent all floats as heap allocated tagged objects or NuN-boxing, and Bigloo also supports NaN-boxing. This provides a flexible environment for comparing these techniques.
In this paper the C back end of the Bigloo and Gambit compilers are used. This way the resulting machine code benefits from the C compiler’s optimizations and in particular excellent architecture specific instruction selection. In both cases the same version of the gcc C compiler is used.
4.1. Overview of the Bigloo Compiler
The Bigloo compiler conforms to the R5RS Scheme specification and adds several extensions, including optional type annotations. Exact rational numbers and complex numbers, which are optional in the R5RS Scheme specification, are not supported.
The Bigloo compiler builds a typed abstract syntax tree from the compiled Scheme program. Several analysis and optimizations refine the initial optional type annotations contained in the source program, the main ones being occurrence-typing (Tobin-Hochstadt and Felleisen, 2010) and storage use analysis (Serrano and Feeley, 1996). During compilation, primitive values have a dual representation: one for polymorphic contexts and one for specific contexts. For instance, when the compiler can establish that a function is always invoked with floats, the compiler assigns the C double type to that function and, when needed, it introduces cast operations from and to the generic representation of floats and double.
The result of this compilation technique is that many local variables and local functions are precisely typed and avoid polymorphic representations. This contributes to removing otherwise necessary heap allocations and casts for floats. The compiler also tracks cases where only floats are stored into vectors, which are then transformed into arrays of C doubles.
Bigloo uses the Boehm-Demers-Weiser mark-and-sweep conservative garbage collector (Boehm and Weiser, 1988). The complexity of collecting dead objects is proportional to the number of allocated objects that are still live, contrary to copying collectors. For such a collector, avoiding allocating short lived floats is crucial for performance.
4.2. Overview of the Gambit Compiler
The Gambit compiler conforms to the R7RS Scheme language. It features the full numerical tower including arbitrary-precision integers, exact rational numbers and complex numbers. Arithmetic operators are generic and the result type depends on the specific values, for example multiplying two complex numbers may yield a number of any type. Like Bigloo, the compiler inlines the small integer and float cases of the dispatch and handles other cases in an out-of-line function.
First-class continuations (call/cc) and tail-calls are fully supported, and stack-overflows are gracefully handled. With the C back end, this requires the use of a trampoline and the management of the stack through an explicit array, stack pointer and stack-overflow checks. This causes a small slowdown when compared to a machine code back end where the trampoline would not be required. Memory management is based on a fast bump-allocator and a stop-and-copy garbage collector.
The compiler front end performs several source-to-source optimizations, such as constant folding, lambda lifting, and function inlining. It does not perform type analysis or type inference. However, the back end uses a simple mechanism to optimize float calculations by tracking the use of their results within basic blocks. The result of a float operation is kept in raw form so that other operations within the same basic block that consume it do not need to unbox the value. A float value is boxed only if it remains live at the end of the basic block.
4.3. Experimental Setup
Self-tagging variants are implemented in Bigloo and Gambit, and tested using a subset of the R7RS benchmarks, which is the standard benchmark suite for Scheme (r7r, 2024). The experiment includes all macro-benchmarks (more than 500 lines of code) and benchmarks whose calculations mainly involve floats (tagged as such in the R7RS benchmarks suite). For the purpose of analyzing results, it is useful to distinguish between benchmarks that use few or no floats and float-intensive benchmarks (fibfp, fft, mbrot, nucleic, pnpoly, ray, simplex, sum1, and sumfp).
Experiments were repeated on the following machines that operate on distinct microarchitectures:
  • Intel - Intel(R) Xeon(R) W-2245, 32GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
  • Amd - AMD Ryzen 7955WX, 62GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
  • M2 - Apple M2 Max, 64GB, macOS 15.4.1, gcc 14.3.0
  • Risc-V - riscv sifive, u74-mc, 8GB, Linux starfive 6.1.31-starfive riscv64, gcc 14.2.0
For each combination of machine, compiler and self-tagging variant, benchmarks are executed 10 times, and each repetition is configured to last at least 5 seconds. Repetitions of each variant and baseline are paired in order of execution to compute relative execution time for each pair. Upcoming sections report geometric means of these relative times, with geometric standard deviations.
To reduce variance on Intel and Amd, benchmarks are executed on the same CPU core by using taskset and address space randomization is disabled with setarch. On M2, no equivalent tools are available and hence the variance is higher on some benchmarks. On Risc-V, the variance remains very low despite not using taskset and setarch.
For measurements that were similar across all machines, a single microarchitectures is shown. Omitted figures can be found in the artifact that comes with this paper (Melançon et al., 2025).
4.4. Memory Profiling and Execution Time
Heap Allocations Relative to Allocated Floats (Bigloo)
Intel(R) Xeon(R) W-2245, 32GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
  self-tagging (1-tag)   self-tagging (2-tag w/ prealloc. zero)   self-tagging (3-tag)
  NaN-boxing   NuN-boxing
Figure 6. Heap allocations of Bigloo with each variant of self-tagging, NaN-boxing, NuN-boxing, and allocated floats. The y-axis shows the number of allocated bytes relative to the version that heap allocates all floats (dashed horizontal red line). 1.00×1.$0$ indicates no change, while $0\times0$ indicates no heap allocation.
Execution Time Relative to NuN-Boxing (Bigloo)
AMD Ryzen 7955WX, 62GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
Intel(R) Xeon(R) W-2245, 32GB, Linux 6.12.31 x86_64, Debian 10.13, gcc 14.2.0
Apple M2 Max, 64GB, macOS 15.4.1, gcc 14.3.0
riscv sifive, u74-mc, 8GB, Linux starfive 6.1.31-starfive riscv64, gcc 14.2.0
  self-tagging (1-tag)   self-tagging (2-tag w/ prealloc. zero)   self-tagging (3-tag)   NaN-boxing